In a go the product of the first four terms is 4 and the second d term is the reciprocal of the fourth term .    
The sum of infinite terms of the gp is?
 
-8,8,-8/3,8/3
All the four are to be proved

Grade:

1 Answers

Arun
5 years ago
 

let 4 nos be a/r2,a/r,ar,ar2.

a= 4

a = 21/2

ar= r/a

r = 1/2

sum of infinite terms of GP = a/1-r = 8 

 

Regards

Arun